The measure of one acute angle of an obtuse triangle is 20 degrees more than the measure of the other acute angle. What is a possible measure of the smaller acute angle?
Answer by Mathtut(3670)

a+a+20+b=180....where a is the small angle and b is the obtuse angle
:
2a+b=160
:
2a=160-b
:
a=80-b/2
: